# Université de Batna-1, Batna 05000, Algeria

### v1

@@ Abstract
QCD, Jets and Jet Substructure, Resummation, Hadron collisions
@@ 1 Introduction
In this paper, we focus on the NLL resummation of the jet mass distribution in Z+jet final states at the LHC proton-proton collisions for jets defined using the trimming algorithm. This groomer, frequently used by the ATLAS experiment, presents unique challenges, particularly in dealing with the emergence of NGLs and CLs. We perform a fixed-order calculation of the trimmed distribution at leading order and compare our results with those obtained from the Monte Carlo program MCFM 9.1 Campbell:2019dru , finding agreement. By subtracting the large logarithms from the MCFM distribution, the resulting constant at small jet mass values is used to extend the accuracy of the perturbative jet mass distribution to NNLL, resumming all logarithms of the form $\alpha_{s}^{n}\ln^{m}(m_{j}/p_{t})$, with $m=\{2n,2n-1,2n-2\}$, in the expanded cumulative distribution.
@@ 2.1 Process kinematics
We examine the process of production of a Z boson and a hard jet in $pp$ collisions at a center-of-mass energy $\sqrt{s}$. To achieve Single Logarithmic (SL) accuracy in resummation, we assume the emitted gluons are strongly ordered in transverse momentum $k_{ti}$, such that at the $n^{\mathrm{th}}$ order: $k_{tn}\ll\cdots\ll k_{t2}\ll k_{t1}\ll p_{t}$, where $p_{t}$ is the jet’s transverse momentum. In this regime, the momenta of the incoming partons $(a)$ and $(b)$, carrying momentum fractions $x_{a}$ and $x_{b}$ of the colliding protons, along with the outgoing hard parton $(j)$ and soft emissions $(i)$, are expressed as:
@@ 2.3 Resummed trimmed jet mass distribution and global form factor
where $\Omega_{\mathcal{B}}$ represents the experimental cuts and $C_{1}(\mathcal{B}_{\delta})$ is the $\mathcal{O}(\alpha_{s})$ correction term, introduced to improve the accuracy up to NNLL. A detailed discussion of this correction term will be provided later.
@@ 3.1 Review of the ungroomed distribution
The form factor for channel $\delta$, in the eikonal approximation (sufficient to capture single and double logarithms), is given at $\mathcal{O}(\alpha_{s})$ for a $k_{t}$, anti-$k_{t}$, or CA-clustered jet of radius $R$ by Dasgupta:2012hg ; Ziani:2021dxr
@@
where the sum is over the dipoles formed by the hard eikonal legs $(i\ell)=\{(ab),(aj),(bj)\}$. The color factors, $\mathcal{C}_{i\ell}$, depend on the nature of the partons involved: $\mathcal{C}_{i\ell}=\mathrm{C_{A}}=\mathrm{N_{c}}$ for dipoles involving gluons and $\mathcal{C}_{i\ell}=2\,\mathrm{C_{F}}-\mathrm{C_{A}}=-1/\mathrm{N_{c}}$ for dipoles involving only quarks or anti-quarks. Here, $\mathrm{C_{A}}=\mathrm{N_{c}}$ and $\mathrm{C_{F}}=(\mathrm{N_{c}}^{2}-1)/(2\mathrm{N_{c}})$ are the QCD color factors with $\mathrm{N_{c}}=3$. The quantity $\xi_{1}=k_{t1}/p_{t}$ is the momentum fraction of the soft emission relative to the jet’s transverse momentum $p_{t}$. The angular antenna function for the emission of the soft gluon $k_{1}$ from the dipole $(i\ell)$ is given by
@@ 3.3 Comparison to Monte Carlo results
To validate our analytical results, we conduct a phenomenological study to compare them with Monte Carlo simulations. Figure 1 displays the differential distribution $\mathrm{d}\sigma/dL$, where $L=\ln(R^{2}/\rho)$, as a function of $L$. The Monte Carlo results are obtained using the fixed-order program MCFM 9.1. We also utilized MadGraph5_aMC@NLO, which produced similar results. Our obtained results are computed using the NNPDF30_nlo parton distribution functions with $\alpha_{s}(M_{Z})=0.118$, interfaced via LHAPDF Buckley:2014ana . The center-of-mass energy is set to $\sqrt{s}=7\,\mathrm{TeV}$, and jets with transverse momentum $p_{t}>150\,\mathrm{GeV}$ and rapidity $y<2.5$ are selected. We set the Z boson decay width to zero to simplify the analysis. The upper plots in figure 1 show the ungroomed distribution for quark (left) and gluon (right) channels, while the lower plots display the trimmed distributions. The jet radius is fixed at $R=0.7$, with substructure parameters $R_{\mathrm{sub}}=0.2$ and $z_{\mathrm{cut}}=0.03$. Renormalization and factorization scales are chosen to be $\mu_{\mathrm{R}}=\mu_{\mathrm{F}}=150.0\,\mathrm{GeV}$.
@@
3.4 NNLL corrections
To achieve NNLL accuracy in the expansion of the cumulative distribution, we must include all terms of the form $\alpha_{s}^{n}\,L^{2n-2}$ at $n^{\mathrm{th}}$ order. These terms can be captured by calculating the fixed-order constant $C_{1}^{(\delta)}$, averaged over the Born configuration $\mathcal{B}_{\delta}$, as defined in ref. Banfi:2010xy
@@ 5 Convolution and numerical results
To illustrate the impact of NGLs and the constant $\alpha_{s}C_{1}$, we plot three configurations of the resummed distribution in figure 5: (a) resumming global logarithms only, (b) resumming both global and non-global logarithms, and (c) the full resummed distribution, valid at NNLL accuracy, including the constant $C_{1}$. The plots show differential distributions in $\sqrt{\rho}$ normalized to the Born cross-section for the global and resummed distributions, and normalized to the NLO cross-section for the distribution accounting for the constant $C_{1}$. Results are shown for both ungroomed and trimmed jets.
Figure 5: Resummed differential distributions of the jet mass observable, comparing trimmed and untrimmed cases. The plots display results for global logarithms, global and non-global logarithms, and the full resummation at NNLL accuracy.
@@
Figure 6 compares our resummed results, valid at NLL accuracy in the exponent and NNLL accuracy in the expansion, to parton shower simulations from Pythia 8.3 Bierlich:2022pfr ; Alwall:2008qv , Herwig++ Bahr:2008pv ; Bellm:2015jjp , and Sherpa 2.2 Sherpa:2019gpd . The Pythia 8 and Herwig++ results are obtained by showering NLO events from MadGraph5_aMC@NLO. Jet finding for both ungroomed and groomed jets is performed using FastJet Cacciari:2011ma and its plugin FastJet-Contrib.
@@ 6 Conclusions
In this paper we have presented a detailed calculation of the resummed trimmed jet mass distribution at NLL accuracy in the exponent and NNLL accuracy in the perturbative expansion of the cumulative distribution. We performed fixed-order calculations at $\mathcal{O}(\alpha_{s})$ for global logarithms and $\mathcal{O}(\alpha_{s}^{2})$ for non-global and clustering logarithms, and presented a resummed result valid to all orders. Our $\mathcal{O}(\alpha_{s})$ results were found to be consistent with Monte Carlo predictions from the MCFM program.
@@
While this work focused on obtaining the NNLL resummed distribution, further steps are necessary to make meaningful comparisons with experimental data, such as that provided by the CMS collaboration CMS:2013kfv . Specifically, the inclusion of NLO corrections via matching to fixed-order calculations is required. Moreover, non-perturbative corrections, including contributions from the underlying event and hadronization, are expected to have a significant impact on the distribution Ziani:2021dxr ; Dasgupta:2012hg , and must be included before any comparison with experimental results. These corrections, along with assessments of statistical and systematic uncertainties, will be addressed in our future work.

### v2

# Batna 05000, Algeria

@@ Abstract
Jets and Jet Substructure, Resummation
@@ 1 Introduction
In this paper, we focus on the NLL resummation of the jet mass distribution in Z+jet final states at the LHC proton-proton collisions for jets defined using the trimming algorithm. This groomer, frequently used by the ATLAS experiment, presents unique challenges, particularly in dealing with the emergence of NGLs and CLs. We perform a fixed-order calculation of the trimmed distribution at leading order and compare our results with those obtained from the Monte Carlo program MCFM 9.1 Campbell:2019dru , finding agreement. By subtracting the large logarithms from the MCFM distribution, the resulting constant at small jet mass values is used to extend the accuracy of the perturbative jet mass distribution to NLL’, accounting for all logarithms of the form $\alpha_{s}^{n}\ln^{m}(m_{j}/p_{t})$, with $m=\{2n,2n-1,2n-2\}$ in the order-by-order expansion of the cumulative distribution Banfi:2010xy .
@@ 2.1 Process kinematics
We examine the process of production of a Z boson and a hard jet in $pp$ collisions at a center-of-mass energy $\sqrt{s}$. The $Z$ plus jet final state has garnered significant interest due to its crucial role in various phenomenological studies. These include its presence as an important background in many processes, its clean experimental signature from Z decays, and its use in probing parton distribution functions and the strong coupling constant. One notable example of such studies is the application of the Transverse Momentum Dependent (TMD) resummation formalism to some observable distributions in this process Sun:2018icb .
To achieve Single Logarithmic (SL) accuracy in resummation, we assume the emitted gluons are strongly ordered in transverse momentum $k_{ti}$, such that at the $n^{\mathrm{th}}$ order: $k_{tn}\ll\cdots\ll k_{t2}\ll k_{t1}\ll p_{t}$, where $p_{t}$ is the jet’s transverse momentum. In this regime, the momenta of the incoming partons $(a)$ and $(b)$, carrying momentum fractions $x_{a}$ and $x_{b}$ of the colliding protons, along with the outgoing hard parton $(j)$ and soft emissions $(i)$, are expressed as:
@@ 2.3 Resummed trimmed jet mass distribution and global form factor
where $\Omega_{\mathcal{B}}$ represents the experimental cuts and $C_{1}(\mathcal{B}_{\delta})$ is the $\mathcal{O}(\alpha_{s})$ correction term, introduced to improve the accuracy up to NLL’. A detailed discussion of this correction term will be provided later.
@@ 3.1 Review of the ungroomed distribution
The form factor for channel $\delta$, in the eikonal approximation (sufficient to capture the soft singularities), is given at $\mathcal{O}(\alpha_{s})$ for a $k_{t}$, anti-$k_{t}$, or CA-clustered jet of radius $R$ by Dasgupta:2012hg ; Ziani:2021dxr
@@
where the sum is over the dipoles formed by the hard/eikonal legs $(i\ell)=\{(ab),(aj),(bj)\}$. The color factors, $\mathcal{C}_{i\ell}$, depend on the nature of the partons involved: $\mathcal{C}_{i\ell}=\mathrm{C_{A}}=\mathrm{N_{c}}$ for dipoles involving gluons and $\mathcal{C}_{i\ell}=2\,\mathrm{C_{F}}-\mathrm{C_{A}}=-1/\mathrm{N_{c}}$ for dipoles involving only quarks or anti-quarks. Here, $\mathrm{C_{A}}=\mathrm{N_{c}}$ and $\mathrm{C_{F}}=(\mathrm{N_{c}}^{2}-1)/(2\mathrm{N_{c}})$ are the QCD color factors with $\mathrm{N_{c}}=3$. The quantity $\xi_{1}=k_{t1}/p_{t}$ is the momentum fraction of the soft emission relative to the jet’s transverse momentum $p_{t}$. The angular antenna function for the emission of the soft gluon $k_{1}$ from the dipole $(i\ell)$ is given by
@@ 3.3 Comparison to Monte Carlo results
To validate our analytical results, we conduct a phenomenological study to compare them with Monte Carlo simulations. Figure 1 displays the analytical differential distribution $\mathrm{d}\sigma/\mathrm{d}L$, obtained by differentiating $\sigma(\rho)$ (using eq. (6) at $\mathcal{O}(\alpha_{s})$) with respect to $L=\ln(R^{2}/\rho)$. This analytical distribution is compared to the Monte Carlo results obtained using the fixed-order program MCFM 9.1. We also utilized MadGraph5_aMC@NLO, which produced similar results. Our obtained results are computed using the NNPDF30_nlo parton distribution functions with $\alpha_{s}(M_{Z})=0.118$, interfaced via LHAPDF Buckley:2014ana . The center-of-mass energy is set to $\sqrt{s}=7\,\mathrm{TeV}$, and jets with transverse momentum $p_{t}>150\,\mathrm{GeV}$ and rapidity $|y|<2.5$ are selected. We set the Z boson decay width to zero to simplify the analysis. The upper plots in figure 1 show the ungroomed distribution for quark (left) and gluon (right) channels, while the lower plots display the trimmed distributions. The jet radius is fixed at $R=0.7$, with substructure parameters $R_{\mathrm{sub}}=0.2$ and $z_{\mathrm{cut}}=0.03$. Renormalization and factorization scales are chosen to be $\mu_{\mathrm{R}}=\mu_{\mathrm{F}}=150.0\,\mathrm{GeV}$.
@@
3.4 NLL’ corrections
To achieve NLL’ accuracy in the expansion of the cumulative distribution, we must include all terms of the form $\alpha_{s}^{n}\,L^{2n-2}$ at $n^{\mathrm{th}}$ order. These terms can be captured by calculating the fixed-order constant $C_{1}^{(\delta)}$, averaged over the Born configuration $\mathcal{B}_{\delta}$, as defined in ref. Banfi:2010xy
@@ 5 Convolution and numerical results
To illustrate the impact of NGLs and the constant $\alpha_{s}C_{1}$, we plot three configurations of the resummed distribution in figure 5: (a) resumming global logarithms only, (b) resumming both global and non-global logarithms, and (c) the full resummed distribution, valid at NLL’ accuracy, including the constant $C_{1}$. The plots show differential distributions in $\sqrt{\rho}$ normalized to the Born cross-section for the global and resummed distributions, and normalized to the NLO cross-section for the distribution accounting for the constant $C_{1}$. Results are shown for both ungroomed and trimmed jets.
Figure 5: Resummed differential distributions of the jet mass observable, comparing trimmed and untrimmed cases. The plots display results for global logarithms, global and non-global logarithms, and the full resummation at NLL’ accuracy.
@@
Figure 6 compares our resummed results, valid at NLL accuracy in the exponent and NLL’ accuracy in the expansion, to parton-level Monte Carlo simulations from Pythia 8.3 Bierlich:2022pfr ; Alwall:2008qv , Herwig++ Bahr:2008pv ; Bellm:2015jjp , and Sherpa 2.2 Sherpa:2019gpd . The Pythia 8 and Herwig++ results are obtained by showering NLO events from MadGraph5_aMC@NLO. Jet finding for both ungroomed and groomed jets is performed using FastJet Cacciari:2011ma and its plugin FastJet-Contrib.
@@ 6 Conclusions
In this paper we have presented a detailed calculation of the resummed trimmed jet mass distribution at NLL accuracy in the exponent and NLL’ accuracy in the perturbative expansion of the cumulative distribution. We performed fixed-order calculations at $\mathcal{O}(\alpha_{s})$ for global logarithms and $\mathcal{O}(\alpha_{s}^{2})$ for non-global and clustering logarithms, and presented a resummed result valid to all orders. Our $\mathcal{O}(\alpha_{s})$ results were found to be consistent with Monte Carlo predictions from the MCFM program.
@@
While this work focused on obtaining the NLL’ resummed distribution, further steps are necessary to make meaningful comparisons with experimental data, such as that provided by the CMS collaboration CMS:2013kfv . Specifically, the inclusion of NLO corrections via matching to fixed-order calculations is required. Moreover, non-perturbative corrections, including contributions from the underlying event and hadronization, are expected to have a significant impact on the distribution Ziani:2021dxr ; Dasgupta:2012hg , and must be included before any comparison with experimental results. These corrections, along with assessments of statistical and systematic uncertainties, will be addressed in our future work.